# MOGO: Residual Quantized Hierarchical Causal Transformer for Real-Time and Infinite-Length 3D Human Motion Generation

### v2

@@ Abstract
I Introduction
Text-to-motion generation is a rapidly evolving research area with growing importance in virtual environments, such as gaming, AR/VR, and humanoid robotics [1]. Recent advances have leveraged large language model (LLM)-based techniques to generate high-quality 3D human motion from text descriptions, typically using vector-quantized variational autoencoders (VQ-VAE) and autoregressive decoding strategies [2, 3, 4, 5]. Meanwhile, diffusion-based approaches have also demonstrated strong generation performance, particularly in terms of motion fidelity and diversity [6].
Despite these achievements, both diffusion- and LLM-based frameworks face practical limitations [7, 8, 9, 10]. Diffusion models often rely on iterative refinement processes, which introduce significant inference latency and hinder their suitability for real-time or interactive applications [7]. On the other hand, LLM-based models, although autoregressive, typically involve large parameter sizes and long context dependencies, leading to high memory and computation costs that challenge deployment on lightweight scenarios [9]. A broader discussion of related research trends is provided in Section II.
@@ II Related Work
Human Motion Generation. Recent advances in human motion generation have enabled conditioning on modalities like text, audio, music, and images [1]. Early deterministic models [11, 12] often produced over-smoothed, unrealistic motions. To address this, stochastic approaches, such as GANs [13, 14] and VAE-based models [15, 16], improved motion diversity. Text-to-motion generation gained prominence with works like [17], which used temporal VAEs to model text-motion distributions. Recently, diffusion-based methods [18, 19, 20, 21, 8, 6] and transformer-based approaches [2, 3, 22] have led the field.
@@
LLM-Based Motion Generation Models. LLM-based architectures have become a cornerstone of text-to-motion generation, leveraging their ability to model sequential data and adapt to varied tasks [2, 3, 4, 5, 22]. Models like MoMask [2] and MMM [4] use masked token modeling to produce high-quality motions but struggle with extendable output and generalization in low-data or out-of-distribution settings due to their bidirectional design. Conversely, autoregressive models like T2M-GPT [3] and AttT2M [5] enable sequential generation, making them suitable for real-time applications and scalable with larger datasets, though they often sacrifice some motion quality. Efforts like MotionGPT [22] integrate multimodal language modeling but face challenges in achieving high-fidelity motion outputs. Our MOGO framework addresses these issues by combining efficient encoding through the MoSA-VQ with a single-pass autoregressive transformer, ensuring both high-quality motion and interactive capabilities.
@@
Hierarchical Transformers. Hierarchical transformer architectures excel in domains like NLP [23, 24], image generation [25], and vision tasks [26, 27]. By processing data at multiple abstraction levels, these models enhance representation capacity and scalability. For instance, Swin Transformers [26] support scalable high-resolution vision, while CogView2 [25] enables high-fidelity image synthesis. In motion generation, hierarchical transformers are underexplored, especially for autoregressive frameworks with residual quantization. Our RQHC-Transformer in MOGO leverages hierarchical modeling to efficiently process multi-layer motion tokens, improving quality and generalization.
@@
Motion Tokenization. Discretizing continuous motion data into tokens v‘ia vector quantization is central to transformer-based motion generation. TM2T [28] introduced VQ-VAE to map motions to discrete sequences. T2M-GPT [3] enhanced token quality with exponential moving average and codebook reset techniques. AttT2M [5] improved quantization through body-part-aware encoding. MoMask [2] advanced this with residual vector quantization (RVQ), producing multi-level tokens for better reconstruction quality. Compared to prior work, our approach introduces learnable feature scaling into the residual vector quantization process, allowing each quantization level to adaptively adjust the magnitude of its residuals. This scaled RVQ mechanism stabilizes the residual distribution across quantization stages, mitigates the risk of later stages collapsing to noise, and ensures more effective codebook utilization. As a result, the model benefits from improved reconstruction fidelity, better token expressiveness, and enhanced training stability.
III Methods
III-A Overview of the MOGO Framework
@@
III-B MoSA-VQ: Motion Scale-Adaptive Residual VQ-VAE
To transform continuous human motion into discrete representations suitable for autoregressive generation, we build upon residual vector quantized variational autoencoders (RVQ-VAEs) and propose MoSA-VQ. While standard RVQ-based approaches [2, 5] hierarchically encode motion features through fixed-stage quantization, they often rely on manually tuned feature ranges and static scaling behavior, which limit their adaptability and quantization efficiency—especially in the presence of diverse and complex motions.
@@ III-B MoSA-VQ: Motion Scale-Adaptive Residual VQ-VAE
$\mathbf{q}^{l}=\mathcal{Q}(\mathbf{W}^{l}\mathbf{r}^{l}+\mathbf{b}^{l}),\quad\bm{\phi}^{l}=(\mathbf{W}^{l})^{-1}(\mathbf{q}^{l}-\mathbf{b}^{l}),$ (1)
where $\mathbf{W}^{l}\in\mathbb{R}^{d\times d}$ is a diagonal scaling matrix, $\mathbf{b}^{l}\in\mathbb{R}^{d}$ is a learnable bias vector, $\mathcal{Q}(\cdot)$ denotes nearest-neighbor quantization via a codebook, and $\bm{\phi}^{l}$ represents the decoded contribution at level $l$.
@@
$\mathbf{r}^{l+1}=\mathbf{r}^{l}-\bm{\phi}^{l},\quad\hat{\mathbf{z}}=\sum_{l=0}^{L}\bm{\phi}^{l}.$ (2)
@@
We adopt an $\ell_{1}$ reconstruction loss to ensure accurate motion recovery, and a standard commitment loss to encourage the encoder residuals $\mathbf{r}^{l}$ to remain close to their quantized counterparts $\bm{\phi}^{l}$. A mild regularization is also applied to the per-layer scaling and bias parameters $(\mathbf{W}^{l},\mathbf{b}^{l})$, following prior work on stable quantization with learnable modulation.
Additionally, we introduce a cross-level decorrelation loss, which explicitly penalizes statistical redundancy between the quantized vector $\bm{\phi}^{l}$ and the subsequent residual $\mathbf{r}^{l+1}$:
$\mathcal{L}_{\text{decor}}=\sum_{l=0}^{L-1}\left\|\mathrm{Cov}\left(\bm{\phi}^{l},\mathbf{r}^{l+1}\right)\right\|_{F}^{2}.$
This loss is motivated by the principle that, in an ideal residual quantization hierarchy, each level should encode novel information orthogonal to what has already been captured. A nonzero covariance between $\bm{\phi}^{l}$ and $\mathbf{r}^{l+1}$ indicates overlapping representation subspaces, leading to inefficient coding and excessive reliance on deeper quantizers. By minimizing their batch-wise covariance, we enforce an approximate statistical orthogonality between levels, which encourages disentangled information flow and hierarchical efficiency.
@@
$\mathcal{L}_{\mathrm{vq}}=\$$\gamma\sum_{l=0}^{L}\underbrace{\left(\left\|\mathbf{W}^{l}\!-\!\mathbf{I}\right\|_{F}^{2}\!+\!\left\|\mathbf{b}^{l}\right\|_{2}^{2}\right)}_{\text{modulation reg.}}\!+\!\beta\sum_{l=0}^{L}\underbrace{\left\|\mathbf{r}^{l}\!-\!\mathrm{sg}[\bm{\phi}^{l}]\right\|_{2}^{2}}_{\text{commitment}}$
$+\lambda\sum_{l=0}^{L-1}\underbrace{\left\|\mathrm{Cov}\left(\bm{\phi}^{l},\mathbf{r}^{l+1}\right)\right\|_{F}^{2}}_{\text{cross-level decorrelation}}+\underbrace{\|\mathbf{m}-\hat{\mathbf{m}}\|_{1}}_{\text{reconstruction}}.$ (3)
@@
III-C RQHC-Transformer: Residual Quantized Hierarchical Causal Transformer
@@
III-D Infinite-Length Generation
@@
III-E Text Condition Alignment for Improved Motion Decoding
@@
TABLE I: Comparison with Motion Generation Models. Bold Boxed = Best, Bold = Second Best, Underline = Third Best.
@@ IV Experiments
IV-A Datasets and Evaluation Metrics
IV-A1 Datasets
We train and evaluate MOGO on two widely used benchmarks, HumanML3D [17] and KIT-ML [31], following the data splits of T2M [17]. To assess out-of-distribution generalization, we further evaluate zero-shot performance on CMP [32], which contains non-daily, combat-style motions. Prompt engineering is disabled during CMP evaluation to ensure fairness.
IV-A2 Evaluation Metrics
We adopt standard metrics from [17]. FID (Frechet Inception Distance) measures motion realism and serves as a primary evaluation metric. R-Precision evaluates text-motion alignment (Top-1, 2, 3). MM-Dist computes the distance between motion and text embeddings. MultiModality measures the variance of motions generated from the same text prompt.
IV-A3 Training Details
@@
IV-B Comparison to State-of-the-art Approaches
@@
TABLE II: Comparison of reconstruction quality between our VAE design and prior motion VAEs.
@@
IV-B1 Reconstruction Quality of Motion Encoder
We compare our VAE design with existing motion VAEs—M2DM, T2M-GPT, MoMask, and MMM—on HumanML3D and KIT-ML (Table II). MOGO achieves the best results on both benchmarks, with an FID of 0.013 on HumanML3D, outperforming MoMask (0.019) and others. On the more challenging KIT-ML dataset, MOGO achieves an FID of 0.037, significantly lower than MoMask (0.112), T2M-GPT (0.472), and MMM (0.641). These results demonstrate the strong reconstruction capability of our VAE across diverse motion distributions.
IV-B2 Quantitative Results.
MOGO demonstrates strong performance across HumanML3D, KIT-ML, and zero-shot CMP benchmarks, particularly on FID and R@3. As showned in Table I, On HumanML3D, our base model achieves an FID of 0.038, outperforming T2M-GPT (0.116), MotionGPT (0.232), and MotionDiffuse (0.630). With TCA, FID slightly increases to 0.064, while R@3 improves from 0.801 to 0.827. On KIT-ML, MOGO attains an FID of 0.191 and R@3 of 0.801, competitive with MotionAnything (FID 0.131, R@3 0.802). Under the zero-shot CMP setting, FID drops to 10.388 (6.873 with TCA), outperforming MotionGPT (15.654) and T2M-GPT (16.092), while R@3 reaches 0.304, over twice that of the best baseline MMM (0.154).
@@ IV-B1 Reconstruction Quality of Motion Encoder
IV-C Ablation Study
IV-C1 Ablation Study on VAE Depth
TABLE III: The impact of different quantization layers on model reconstruction quality when the codebook size is $8192\times 128$. Bold face indicates the best result.
@@ IV-C1 Ablation Study on VAE Depth
We evaluate the impact of quantization depth by varying the number of layers while fixing the codebook size to $8192\times 128$ (Table VIII). A single-layer quantizer performs poorly (FID: 0.070), indicating limited capacity. Increasing to 3 layers significantly improves performance (FID: 0.021), and 6 layers yields the best overall results (FID: 0.016, R@1: 0.510, MM-Dist: 2.989). Adding a 7th layer offers no further gain and slightly worsens MM-Dist, suggesting diminishing returns. Overall, 6-layer quantization provides the best balance between expressiveness and generalization.
IV-C2 Ablation Study on Layer-Head Configuration
TABLE IV: Impact of model depths and head configurations on HumanML3D results (codebook size: $8192\times 128$). Bold indicates the best performance.
@@ IV-C2 Ablation Study on Layer-Head Configuration
We study the impact of Transformer depth and attention head configurations on HumanML3D (Table IV). Starting from a uniform setup ([4, 4, 4, 4, 4, 4] layers, [6, 6, 6, 6, 6, 6] heads), the model yields suboptimal results (FID 0.101, MM-Dist 3.359). As we adopt deeper, hierarchically structured layers and progressively narrower heads, performance improves consistently. The best configuration ([18, 16, 8, 4, 2, 2] layers, [16, 12, 6, 2, 2, 2] heads) achieves the lowest FID (0.079), highest R@1 (0.505), and lowest MM-Dist (3.002).

### v3

@@ Abstract
Codehttps://github.com/MiRECoFu/Mogo
Introduction
Text-to-motion generation is a rapidly evolving research area with growing importance in virtual environments, such as gaming, AR/VR, and humanoid robotics (Zhu et al. 2023). Recent advances have leveraged large language model (LLM)-based techniques to generate high-quality 3D human motion from text descriptions, typically using vector-quantized variational autoencoders (VQ-VAE) and autoregressive decoding strategies (Guo et al. 2024; Zhang et al. 2023a; Pinyoanuntapong et al. 2024; Zhong et al. 2023). Meanwhile, diffusion-based approaches (Sun et al. 2025; Wang et al. 2025) have also demonstrated strong generation performance, particularly in terms of motion fidelity and diversity (Tevet et al. 2022).
Despite these achievements, both diffusion- and LLM-based frameworks face practical limitations (Rombach et al. 2022; Chen et al. 2023; Brown et al. 2020; Raffel et al. 2020). Diffusion models often rely on iterative refinement processes, which introduce significant inference latency and hinder their suitability for real-time or interactive applications (Rombach et al. 2022). On the other hand, LLM-based models, although autoregressive, typically involve large parameter sizes and long context dependencies, leading to high memory and computation costs that challenge deployment on lightweight scenarios (Brown et al. 2020). A broader discussion of related research trends is provided in Section Related Work.
@@ Introduction
Human Motion Generation. Recent advances in human motion generation have enabled conditioning on modalities like text, audio, music, and images (Zhu et al. 2023). Early deterministic models (Ahuja and Morency 2019; Ghosh et al. 2021) often produced over-smoothed, unrealistic motions. To address this, stochastic approaches, such as GANs (Cai et al. 2018; Wang et al. 2020) and VAE-based models (Guo et al. 2022c; Petrovich et al. 2021), improved motion diversity. Text-to-motion generation gained prominence with works like (Guo et al. 2022a; Zhou et al. 2024; Lu et al. 2025; Fan et al. 2025), which used temporal VAEs to model text-motion distributions. Recently, diffusion-based methods (Kim et al. 2023; Zhang et al. 2024, 2023b; Kong et al. 2023; Chen et al. 2023; Tevet et al. 2022) and transformer-based approaches (Guo et al. 2024; Zhang et al. 2023a; Jiang et al. 2023) have led the field.
@@
LLM-Based Motion Generation Models. LLM-based architectures have become a cornerstone of text-to-motion generation, leveraging their ability to model sequential data and adapt to varied tasks (Guo et al. 2024; Zhang et al. 2023a; Pinyoanuntapong et al. 2024; Zhong et al. 2023; Jiang et al. 2023). Models like MoMask (Guo et al. 2024) and MMM (Pinyoanuntapong et al. 2024) use masked token modeling to produce high-quality motions but struggle with extendable output and generalization in low-data or out-of-distribution settings due to their bidirectional design. Conversely, autoregressive models like T2M-GPT (Zhang et al. 2023a) and AttT2M (Zhong et al. 2023) enable sequential generation, making them suitable for real-time applications and scalable with larger datasets, though they often sacrifice some motion quality. Efforts like MotionGPT (Jiang et al. 2023) integrate multimodal language modeling but face challenges in achieving high-fidelity motion outputs. Our MOGO framework addresses these issues by combining efficient encoding through the MoSA-VQ with a single-pass autoregressive transformer, ensuring both high-quality motion and interactive capabilities.
@@
Hierarchical Transformers. Hierarchical transformer architectures excel in domains like NLP (Nawrot et al. 2022; Pappagari et al. 2019; Fang et al. 2025), image generation (Ding et al. 2022), and vision tasks (Liu et al. 2021; Chen et al. 2022, 2024, 2025). By processing data at multiple abstraction levels, these models enhance representation capacity and scalability. For instance, Swin Transformers (Liu et al. 2021) support scalable high-resolution vision, while CogView2 (Ding et al. 2022) enables high-fidelity image synthesis. In motion generation, hierarchical transformers are underexplored, especially for autoregressive frameworks with residual quantization. Our RQHC-Transformer in MOGO leverages hierarchical modeling to efficiently process multi-layer motion tokens, improving quality and generalization.
@@
Motion Tokenization. Discretizing continuous motion data into tokens v‘ia vector quantization is central to transformer-based motion generation. TM2T (Guo et al. 2022b) introduced VQ-VAE to map motions to discrete sequences. T2M-GPT (Zhang et al. 2023a) enhanced token quality with exponential moving average and codebook reset techniques. AttT2M (Zhong et al. 2023) improved quantization through body-part-aware encoding. MoMask (Guo et al. 2024) advanced this with residual vector quantization (RVQ), producing multi-level tokens for better reconstruction quality. Compared to prior work, our approach introduces learnable feature scaling into the residual vector quantization process, allowing each quantization level to adaptively adjust the magnitude of its residuals. This scaled RVQ mechanism stabilizes the residual distribution across quantization stages, mitigates the risk of later stages collapsing to noise, and ensures more effective codebook utilization. As a result, the model benefits from improved reconstruction fidelity, better token expressiveness, and enhanced training stability.
Methods
Overview of the MOGO Framework
@@ Methods
MoSA-VQ: Motion Scale-Adaptive Residual VQ-VAE
To transform continuous human motion into discrete representations suitable for autoregressive generation, we build upon residual vector quantized variational autoencoders (RVQ-VAEs) and propose MoSA-VQ. While standard RVQ-based approaches (Guo et al. 2024; Zhong et al. 2023) hierarchically encode motion features through fixed-stage quantization, they often rely on manually tuned feature ranges and static scaling behavior, which limit their adaptability and quantization efficiency—especially in the presence of diverse and complex motions.
@@
$\mathbf{q}^{l}=\mathcal{Q}(\mathbf{W}^{l}\mathbf{r}^{l}+\mathbf{b}^{l}),\quad\boldsymbol{\phi}^{l}=(\mathbf{W}^{l})^{-1}(\mathbf{q}^{l}-\mathbf{b}^{l}),$ (1)
where $\mathbf{W}^{l}\in\mathbb{R}^{d\times d}$ is a diagonal scaling matrix, $\mathbf{b}^{l}\in\mathbb{R}^{d}$ is a learnable bias vector, $\mathcal{Q}(\cdot)$ denotes nearest-neighbor quantization via a codebook, and $\boldsymbol{\phi}^{l}$ represents the decoded contribution at level $l$.
@@
$\mathbf{r}^{l+1}=\mathbf{r}^{l}-\boldsymbol{\phi}^{l},\quad\hat{\mathbf{z}}=\sum_{l=0}^{L}\boldsymbol{\phi}^{l}.$ (2)
@@
We adopt an $\ell_{1}$ reconstruction loss to ensure accurate motion recovery, and a standard commitment loss to encourage the encoder residuals $\mathbf{r}^{l}$ to remain close to their quantized counterparts $\boldsymbol{\phi}^{l}$. A mild regularization is also applied to the per-layer scaling and bias parameters $(\mathbf{W}^{l},\mathbf{b}^{l})$, following prior work on stable quantization with learnable modulation.
Additionally, we introduce a cross-level decorrelation loss, which explicitly penalizes statistical redundancy between the quantized vector $\boldsymbol{\phi}^{l}$ and the subsequent residual $\mathbf{r}^{l+1}$:
$\mathcal{L}_{\text{decor}}=\sum_{l=0}^{L-1}\left\|\mathrm{Cov}\left(\boldsymbol{\phi}^{l},\mathbf{r}^{l+1}\right)\right\|_{F}^{2}.$
This loss is motivated by the principle that, in an ideal residual quantization hierarchy, each level should encode novel information orthogonal to what has already been captured. A nonzero covariance between $\boldsymbol{\phi}^{l}$ and $\mathbf{r}^{l+1}$ indicates overlapping representation subspaces, leading to inefficient coding and excessive reliance on deeper quantizers. By minimizing their batch-wise covariance, we enforce an approximate statistical orthogonality between levels, which encourages disentangled information flow and hierarchical efficiency.
@@
$\mathcal{L}_{\mathrm{vq}}=\$$\gamma\sum_{l=0}^{L}\underbrace{\left(\left\|\mathbf{W}^{l}\!-\!\mathbf{I}\right\|_{F}^{2}\!+\!\left\|\mathbf{b}^{l}\right\|_{2}^{2}\right)}_{\text{modulation reg.}}\!+\!\beta\sum_{l=0}^{L}\underbrace{\left\|\mathbf{r}^{l}\!-\!\mathrm{sg}[\boldsymbol{\phi}^{l}]\right\|_{2}^{2}}_{\text{commitment}}$
$+\lambda\sum_{l=0}^{L-1}\underbrace{\left\|\mathrm{Cov}\left(\boldsymbol{\phi}^{l},\mathbf{r}^{l+1}\right)\right\|_{F}^{2}}_{\text{cross-level decorrelation}}+\underbrace{\|\mathbf{m}-\hat{\mathbf{m}}\|_{1}}_{\text{reconstruction}}.$ (3)
@@
RQHC-Transformer: Residual Quantized Hierarchical Causal Transformer
@@
Infinite-Length Generation
@@
Text Condition Alignment for Improved Motion Decoding
@@
Table 1: Comparison with motion generation models. Bold boxed entries indicate the best performance, bold entries indicate the second-best, and underlined entries indicate the third-best.
@@
Datasets and Evaluation Metrics
Datasets
We train and evaluate MOGO on two widely used benchmarks, HumanML3D (Guo et al. 2022a) and KIT-ML (Plappert et al. 2016), following the data splits of T2M (Guo et al. 2022a). To assess out-of-distribution generalization, we further evaluate zero-shot performance on CMP (Liao et al. 2024), which contains non-daily, combat-style motions. Prompt engineering is disabled during CMP evaluation to ensure fairness.
Evaluation Metrics
We adopt standard metrics from (Guo et al. 2022a). FID (Frechet Inception Distance) measures motion realism and serves as a primary evaluation metric. R-Precision evaluates text-motion alignment (Top-1, 2, 3). MM-Dist computes the distance between motion and text embeddings. MultiModality measures the variance of motions generated from the same text prompt.
Training Details
@@
Comparison to State-of-the-art Approaches
@@
Table 2: Comparison of reconstruction quality between our VAE design and prior motion VAEs.
Reconstruction Quality of Motion Encoder
We compare our VAE design with existing motion VAEs—M2DM, T2M-GPT, MoMask, and MMM—on HumanML3D and KIT-ML (Table 2). MOGO achieves the best results on both benchmarks, with an FID of 0.013 on HumanML3D, outperforming MoMask (0.019) and others. On the more challenging KIT-ML dataset, MOGO achieves an FID of 0.037, significantly lower than MoMask (0.112), T2M-GPT (0.472), and MMM (0.641). These results demonstrate the strong reconstruction capability of our VAE across diverse motion distributions.
Quantitative Results.
MOGO demonstrates strong performance across HumanML3D, KIT-ML, and zero-shot CMP benchmarks, particularly on FID and R@3. As showned in Table 1, On HumanML3D, our base model achieves an FID of 0.064, outperforming T2M-GPT (0.116), MotionGPT (0.232), and MotionDiffuse (0.630). With TCA, FID slightly improve to 0.038, while R@3 improves from 0.801 to 0.827. On KIT-ML, MOGO attains an FID of 0.191 and R@3 of 0.801, competitive with MotionAnything (FID 0.131, R@3 0.802). Under the zero-shot CMP setting, FID drops to 10.388 (6.873 with TCA), outperforming MotionGPT (15.654) and T2M-GPT (16.092), while R@3 reaches 0.304, over twice that of the best baseline MMM (0.154).
@@
Ablation Study
Ablation Study on VAE Depth
@@
Table 3: The impact of different quantization layers on model reconstruction quality when the codebook size is $8192\times 128$. Bold face indicates the best result.
We evaluate the impact of quantization depth by varying the number of layers while fixing the codebook size to $8192\times 128$ (Table 3). A single-layer quantizer performs poorly (FID: 0.070), indicating limited capacity. Increasing to 3 layers significantly improves performance (FID: 0.021), and 6 layers yields the best overall results (FID: 0.016, R@1: 0.510, MM-Dist: 2.989). Adding a 7th layer offers no further gain and slightly worsens MM-Dist, suggesting diminishing returns. Overall, 6-layer quantization provides the best balance between expressiveness and generalization.
Ablation Study on Layer-Head Configuration
@@
Table 4: Impact of model depths and head configurations on HumanML3D results (codebook size: $8192\times 128$). Bold indicates the best performance.
We study the impact of Transformer depth and attention head configurations on HumanML3D (Table 4). Starting from a uniform setup ([4, 4, 4, 4, 4, 4] layers, [6, 6, 6, 6, 6, 6] heads), the model yields suboptimal results (FID 0.101, MM-Dist 3.359). As we adopt deeper, hierarchically structured layers and progressively narrower heads, performance improves consistently. The best configuration ([18, 16, 8, 4, 2, 2] layers, [16, 12, 6, 2, 2, 2] heads) achieves the lowest FID (0.079), highest R@1 (0.505), and lowest MM-Dist (3.002).